2 kg mass compresses a spring 1.2 m. If the object has 160 J of energy, calculate the spring constant of the spring.
Answer #1

   by the energy conservation , energy of object is changed in spring motion

so using the formula   U= 0.5*k*x^2       here U=energy=160 J; x= compression= 1.2 m; k= spring constant

using the values

                       160= 0.5*k*1.2*1.2

                        k= 222.22 N/m